# Overcoming the Stability Gap in Continual Learning

### v3

@@ 5.1 Experimental Setup
Continual Learning Procedure. We aim to study CL in an industry-like setting, requiring a large-scale data stream with numerous object categories. However, it is difficult to find a suitable dataset that is well-curated and suitable for large-scale CL experiments. Therefore, we construct a large-scale data stream by combining ImageNet-1K with another dataset, Places365-LT which is a variant of the Places365 dataset. Places365 is challenging (Liu et al., 2019) and is widely used for out-of-distribution (OOD) detection with ImageNet pre-trained models (Zhu et al., 2022).
@@
Network Architecture. We choose network architecture that is amenable to LoRA and performs well in offline training with ImageNet-1K compared to similar-sized DNNs. In our main results, we study CL using the ConvNeXtV2-Femto (Woo et al., 2023) CNN that has been pre-trained on ImageNet-1K using a fully convolutional masked autoencoder framework followed by supervised fine-tuning on ImageNet-1K. While ResNet18 is widely used in CL, it under-performs other lightweight CNNs in CL (Hayes & Kanan, 2022; Harun et al., 2023b). ConvNeXtV2-Femto has $5.2$M parameters, which is 2×2$\times$2 × less than ResNet18’s $11.6$M parameters, and it outperforms ResNet18 by absolute 8.47% in final top-1 accuracy on ImageNet-1K.
It is worth noting that LoRA is not appropriate for CNN architectures (e.g., ResNet) that lack $1\times 1$ convolutional layers, but it can be used with ViT, ConvNeXtV1, ConvNeXtV2, and other DNN architectures with these linear layers. Each block of ConvNeXt consists of one 2D convolutional layer and two $1\times 1$ convolutional layers. For experiments using LoRA, the $1\times 1$ convolutional layers in ConvNeXt blocks are modified to incorporate LoRA’s weights using Equation 6. The number of trainable LoRA weights is $0.92$M, which is much less than the total number of hidden layer parameters ($5.08$M). Based on prior work, early layers in the network are universal feature extractors and are little altered during CL (Ramasesh et al., 2021a; Ebrahimi et al., 2020; Pellegrini et al., 2020), so we freeze the first 4 blocks of the CNN in all experiments, leaving the remaining 8 blocks plastic (97.7% of the parameters) which consist of 5.08M parameters.
@@ 5.2.1 Evaluating Our Hypotheses
Baselines. In our experiments, we evaluate each of the methods in Sec. 4 individually and the combined SGM method. As baselines, we compare SGM with rehearsal against vanilla rehearsal, which uses unlimited storage for rehearsal without any additional components, as well as joint models (upper bound). The joint models are jointly trained on ImageNet and the CL batches seen up to the current batch. We also include naive finetune which is a vanilla variant without rehearsal and serves as a lower bound. Finally, we compare with the output layer only which updates the output layer with rehearsal while freezing all other layers.
@@
Hypothesis 2. Our second hypothesis was that the stability gap in CIL is caused by excessive network plasticity, which we tested by using LoRA and OOCF. Fig. 2(c) shows the average results on ImageNet-1K across the 5 rehearsal sessions for LoRA vs. when only the output layer, top 8 blocks (vanilla), or entire network are trainable. This reveals that plasticity plays a major role in the stability gap; however, this does not translate directly into the number of trainable parameters since LoRA includes the output layer but exhibits a smaller decrease in performance than training only the output layer. Unlike others, LoRA fully recovers the old performance. As seen in Table 1, both OOCF and especially LoRA reduce the stability gap.
@@
Table 2: Average CIL Results. Results after learning ImageNet-1K followed by Places365-LT over 5 rehearsal sessions. $\mu$ denotes average accuracy and $\alpha$ is final accuracy on 1365 classes. $\sigma$ stands for final accuracy on ImageNet-1K only. We report the average over $6$ data orderings with standard deviation ($\pm$).
@@
CIL with Multiple Data Orderings. To ensure the robustness of our findings, we extend our experiments to multiple data orderings. In addition to studying SGM and vanilla methods, we also examine training only the output layer during rehearsal. The averaged results across 6 data orderings are given in Table 2. We find that SGM consistently mitigates the stability gap. Compared to vanilla, SGM provides 20×20$\times$20 × more stability, $\times4$ more plasticity, and 15.5×15.5$\times$15.5 × more continual knowledge. SGM also outperforms training only the output layer in all criteria. This indicates that updating representations in hidden layers besides the output layer using SGM is critical for learning new knowledge and retaining old knowledge. Moreover, SGM also maintains a balance between stability and plasticity (see Fig. 6).
Figure 3: Speed of acquiring new knowledge. SGM needs fewer updates and TFLOPs than vanilla to reach 99% of the best accuracy on new classes (highlighted).
@@ 5.2.2 Learning Efficiency
One of our goals in studying CL and the stability gap is to enable more computationally efficient training. To study if SGM achieves this goal, we evaluated performance during CIL, where we measured performance on old and new classes every 10 iterations. For new classes, we measured the number of updates and FLOPs (floating-point operations) needed to achieve 99% of the best accuracy.
As shown in Fig. 3, SGM learns new classes much faster than vanilla rehearsal, and on average it is 61.8% more efficient in terms of network updates. Moreover, because the curve shows a decreasing trend as learning progresses this means that SGM becomes a more efficient learner over time, with it generally requiring fewer updates and TFLOPs for the current batch than previous batches. For old class performance, we measured the number of iterations during rehearsal needed to recover the performance of a joint model (upper bound) on ImageNet-1K for vanilla and SGM. As shown in Table 3, SGM requires far fewer iterations to recover old performance whereas vanilla rehearsal (without SGM) fails to fully recover old performance. Compared to a joint model, SGM provides a 16.7×16.7$\times$16.7 × speedup in number of network updates and a 31.9×31.9$\times$31.9 × speedup in TFLOPs (see Fig. 5). Moreover, SGM requires 18×18$\times$18 × less training time than the joint model on the same hardware.
@@ 6.1 IID Continual Learning
Table 4: IID CL Results. Results after learning ImageNet-1K followed by Places365-LT over 5 rehearsal sessions. $\mu$ denotes average accuracy and $\alpha$ is final accuracy on 1365 classes. $\sigma$ stands for final accuracy on ImageNet-1K only. We report the average over $5$ data orderings with standard deviation ($\pm$).
@@ 6.2 Storage Constrained Offline Continual Learning
After being pre-trained on ImageNet-1K, a model sequentially learns 5 incremental batches of data (73 classes per batch) from Places365-Standard in offline CL with CIL ordering. Each rehearsal is compute-constrained where the model is updated over 1200 minibatches. Each minibatch consists of 256 samples where 50% are selected randomly from the current CL batch and 50% from data seen in earlier CL batches and ImageNet-1K. During learning a new batch, the model rehearses old data from storage which is bounded by a maximum number of samples e.g., 192K and 24K corresponding to $6.4\%$ and $0.8\%$ of the entire dataset (ImageNet and Places combined) respectively. As shown in Table 5, SGM improves each method’s performance in all criteria. When the storage is bounded by 24K samples, SGM improves final accuracy by absolute $11.24\%$ (DERpp) and $14.08\%$ (GDumb) and provides 2.4×2.4$\times$2.4 × and 3.1×3.1$\times$3.1 × more stability for DERpp and GDumb, respectively.
@@ 6.3 Storage Constrained Online Continual Learning
We also assess SGM’s efficacy in an online CL setting using a state-of-the-art online CL method, REMIND (Hayes et al., 2020). We conduct storage-constrained CL experiments with CIL data ordering, where we combine SGM with REMIND while using identical configurations. Here we use the same SSL pre-trained ConvNeXtV2-Femto. Since performing online CL experiments on large-scale datasets is computationally expensive, we choose a small dataset, CUB-200 as the 2nd dataset after ImageNet-1K. CUB-200 is more challenging than other small datasets (Lee et al., 2023). After being pre-trained on ImageNet-1K, a model learns CUB-200 in sample-by-sample manner i.e., after receiving a new sample, it takes one SGD step using 51 samples (50 old samples and 1 new sample). The storage is constrained by a maximum number of samples e.g., 80K and 20K corresponding to $6.2\%$ and $1.5\%$ of the entire dataset (ImageNet and CUB combined) respectively. As shown in Table 6, SGM combined with REMIND outperforms standalone REMIND by large margins in all metrics. When the storage is bounded by 20K samples, SGM improves REMIND’s final accuracy by absolute $8.64\%$ and reduces REMIND’s stability gap by a factor of 3.95×3.95$\times$3.95 ×.
@@ 6.4 Supporting Studies
Class-Balanced Rehearsal. Since our main experiments are based on conventional rehearsal without class balance, we also conduct experiments where class-balanced rehearsal is used and improves performance over prior results. Compared to vanilla, SGM reduces the stability gap by a factor of 7.3×7.3$\times$7.3 ×, when both vanilla and SGM use class-balanced rehearsal. SGM also achieves continual knowledge transfer (see Appendix D.2).
Supervised Pre-Training. In the main results, SGM reduces the stability gap using the self-supervised pre-trained ConvNeXtV2 model, and self-supervised models are known to perform better in CL (Gallardo et al., 2021). To examine how SGM performs without a self-supervised backbone, we also conduct experiments with the supervised pre-trained ConvNeXtV1 (Liu et al., 2022). We find that SGM mitigates the stability gap without the self-supervised backbone. Compared to vanilla, SGM achieves 6×6$\times$6 × more stability, 3.9×3.9$\times$3.9 × more plasticity, and 35×35$\times$35 × more continual knowledge, when both of them use ConvNeXtV1 (see Appendix D.3).
Vision Transformers. We also study the behavior of SGM with ViTs, which now rival CNNs. Compared to vanilla, SGM reduces the stability gap by 2.4×2.4$\times$2.4 ×, when both vanilla and SGM use ViT (see Appendix D.4).
@@ 7 Limitations & Future Work
Our work is built upon the assumption that a well-trained DNN is provided and performance on the original tasks must be maintained or improved. If this assumption is violated, as seen in some CL papers that do not start with pre-trained models, then SGM’s dependence on LoRA would impair its utility. Although most of the components of SGM, such as weight initialization, dynamic soft targets, and OOCF, can be readily applied to DNNs trained from scratch, LoRA cannot be used because DNN layers without LoRA would remain frozen and unlearned. Applying SGM in this setting would require a variant of LoRA where LoRA could be used for all layers from the start of training such that the DNN rivaled a jointly trained model without LoRA. In this work, we used pre-trained ConvNeXtV1, ConvNeXtV2, and MobileViT DNNs, which are amenable to LoRA. We considered using LoRA to limit network plasticity. Future work could investigate different LoRA variants and network adapters (Han et al., 2024) for SGM to mitigate the stability gap.
Our study focused on image classification tasks, while regression tasks are also critical in industry. Future work could explore SGM in regression tasks. Furthermore, it would be interesting to study SGM in other domains, such as object detection (Acharya et al., 2020) and semantic segmentation (Zhang et al., 2022). CL methods developed only on small datasets often do not scale to larger datasets such as ImageNet-1K (Zhou et al., 2023), so we focused on a combination of ImageNet-1K and Places365. Due to computational limitations and a lack of readily available large-scale image classification datasets, we could not study CL for more than 1365 classes. In future work, assessing how well SGM scales with increasing dataset sizes and class counts would be valuable. We hope our work will inspire future research in finding better approaches to prevent model decay due to concept drift.
@@ 8 Conclusion
Although the CL community has shown remarkable progress in mitigating catastrophic forgetting, almost all CL methods remain unsuited for practical applications such as addressing model decay (Verwimp et al., 2024; Harun et al., 2023a). To eliminate the need for frequent retraining of deployed DNNs, CL has to combat model decay, mitigate the stability gap, and be more computationally efficient than retraining from scratch in large-scale CL settings. Focusing on this scenario, we showed that SGM meets these criteria and that it is a candidate solution for addressing model decay using CL. With the growing energy usage of deep learning models (Luccioni et al., 2022; Patterson et al., 2021; Wu et al., 2022), reducing the need for retraining from scratch could significantly contribute to reducing carbon emissions associated with training DNNs. Our work is among the pioneering efforts in CL that align with these real-world challenges, and we hope it encourages the CL community to focus on this critical issue.
@@ Acknowledgments
This work was supported in part by NSF awards #1909696, #2326491, and #2125362. The views and conclusions contained herein are those of the authors and should not be interpreted as representing the official policies or endorsements of any sponsor. We thank Jhair Gallardo, Junyu Chen, Shikhar Srivastava, and Robik Shrestha for their feedback and comments on the manuscript.
@@ Appendix C Additional Implementation Details
Storage Constrained Offline CL with DERpp and GDumb. We describe settings used in Sec. 6.2 where we combine SGM with DERpp and GDumb while using identical settings e.g., the same ImageNet-1K pre-trained ConvNeXt V2 Femto network and the same optimizer settings. Each model pre-trained on ImageNet-1K learns Places365-Standard in 5 batches subsequently (73 categories per batch). Each rehearsal session performs a total of $1200$ iterations with $256$ samples per iteration. DERpp employs distillation and regularization along with rehearsal to prevent catastrophic forgetting. It regularizes loss on old samples and uses an additional distillation loss on logits of old samples for promoting consistency. We set coefficients $\alpha=0.1$ and $\beta=0.9$ for distillation and regularization respectively. GDumb randomly removes a sample from the largest class when the buffer reaches its maximum capacity and maintains a class-balanced buffer. For all methods, the buffer is bounded by a maximum number of samples ($80\%$ ImageNet-1K + $20\%$ Places365-Standard). DERpp, GDumb, and SGM use an initial learning rate of $1\times 10^{-3}$, $1\times 10^{-3}$, and $1.5\times 10^{-3}$, respectively, for batch size $256$. The joint model (upper bound) uses an initial learning rate of $10^{-2}$ and $12$K iterations with $256$ samples per iteration. We assess performance during rehearsal every 100 minibatches to compute the metrics.

### v4

@@ 5.1 Experimental Setup
Continual Learning Procedure. We aim to study CL in an industry-like setting, requiring a large-scale data stream with numerous object categories. However, it is difficult to find a suitable dataset well-curated and suitable for large-scale CL experiments. Therefore, we construct a large-scale data stream by combining ImageNet-1K with another dataset, Places365-LT, which is a variant of the Places365 dataset. Places365 is challenging (Liu et al., 2019) and is widely used for out-of-distribution (OOD) detection with ImageNet pre-trained models (Zhu et al., 2022).
@@
Network Architecture. We choose a network architecture that is amenable to LoRA and performs well in offline training with ImageNet-1K compared to similar-sized DNNs. In our main results, we study CL using the ConvNeXtV2-Femto (Woo et al., 2023) CNN that has been pre-trained on ImageNet-1K using a fully convolutional masked autoencoder framework followed by supervised fine-tuning on ImageNet-1K. While ResNet18 is widely used in CL, it under-performs other lightweight CNNs in CL (Hayes & Kanan, 2022; Harun et al., 2023b). ConvNeXtV2-Femto has $5.2$M parameters, which is 2×2$\times$2 × less than ResNet18’s $11.6$M parameters, and it outperforms ResNet18 by absolute 8.47% in final top-1 accuracy on ImageNet-1K.
It is worth noting that LoRA is not appropriate for CNN architectures (e.g., ResNet) that lack $1\times 1$ convolutional layers, but it can be used with ViT, ConvNeXtV1, ConvNeXtV2, and other DNN architectures with these linear layers. Each block of ConvNeXt consists of one 2D convolutional layer and two $1\times 1$ convolutional layers. For experiments using LoRA, the $1\times 1$ convolutional layers in ConvNeXt blocks are modified to incorporate LoRA’s weights using Equation 6. The number of trainable LoRA weights is $0.92$M, which is much less than the total number of hidden layer parameters ($5.08$M). Based on prior work, early layers in the network are universal feature extractors and are little altered during CL (Ramasesh et al., 2021a; Ebrahimi et al., 2020; Pellegrini et al., 2020; Harun et al., 2024), so we freeze the first 4 blocks of the CNN in all experiments, leaving the remaining 8 blocks plastic (97.7% of the parameters) which consist of 5.08M parameters.
@@ 5.2.1 Evaluating Our Hypotheses
Baselines. In our experiments, we evaluate each of the methods in Sec. 4 individually and the combined SGM method. As baselines, we compare SGM with rehearsal against vanilla rehearsal, which uses unlimited storage for rehearsal without any additional components, as well as joint models (upper bound). The joint models are jointly trained on ImageNet and the CL batches seen up to the current batch. We also include naive finetuning which is a vanilla variant without rehearsal and serves as a lower bound. Finally, we compare with the output layer only which updates the output layer with rehearsal while freezing all other layers.
@@
Hypothesis 2. Our second hypothesis was that the stability gap in CIL is caused by excessive network plasticity, which we tested by using LoRA and OOCF. Fig. 2(c) shows the average results on ImageNet-1K across the 5 rehearsal sessions for LoRA vs. when only the output layer, top 8 blocks (vanilla), or entire network are trainable. This reveals that plasticity plays a major role in the stability gap; however, this does not translate directly into the number of trainable parameters since LoRA includes the output layer but exhibits a smaller decrease in performance than training only the output layer. Unlike others, LoRA fully recovers old performance. As seen in Table 1, both OOCF and especially LoRA reduce the stability gap.
@@
Table 2: Average CIL Results. Results after learning ImageNet-1K followed by Places365-LT over 5 rehearsal sessions. $\mu$ denotes average accuracy (%) over rehearsal sessions and $\alpha$ is final accuracy (%) on all 1365 classes. $\sigma$ stands for final accuracy (%) on ImageNet-1K only. #$\#P$ denotes the trainable parameters in Millions. We report the average over $6$ data orderings with standard deviation ($\pm$).
@@
CIL with Multiple Data Orderings. To ensure the robustness of our findings, we extend our experiments to multiple data orderings. In addition to studying SGM and vanilla rehearsal, we examine only training the output layer during rehearsal. The averaged results across 6 data orderings are given in Table 2. We find that SGM consistently mitigates the stability gap. Compared to vanilla, SGM provides 20×20$\times$20 × more stability, $\times4$ more plasticity, and 15.5×15.5$\times$15.5 × more continual knowledge. SGM also outperforms training only the output layer in all criteria. This indicates that updating representations in hidden layers besides the output layer using SGM is critical for learning new and retaining old knowledge. Moreover, SGM balances stability and plasticity (see Fig. 6).
Figure 3: Speed of acquiring new knowledge. SGM requires fewer updates and TFLOPs than vanilla to reach 99% of the best accuracy on new classes (highlighted).
@@ 5.2.2 Learning Efficiency
One of our goals in studying CL and the stability gap is to enable more computationally efficient training. To study whether SGM achieves this goal, we evaluated performance during CIL, where we measured performance on old and new classes every 10 iterations. For new classes, we measured the number of updates and FLOPs (floating-point operations) needed to achieve 99% of the best accuracy.
As shown in Fig. 3, SGM learns new classes much faster than vanilla rehearsal, and on average, it is 61.8% more efficient in terms of network updates. Moreover, because the curve shows a decreasing trend as learning progresses this means that SGM becomes a more efficient learner over time, with fewer updates and TFLOPs for the current batch than previous batches being needed. For old class performance, we measured the number of iterations during rehearsal required to recover the performance of a joint model (upper bound) on ImageNet-1K for vanilla and SGM. As shown in Table 3, SGM requires far fewer iterations to recover old performance whereas vanilla rehearsal (without SGM) fails to recover old performance fully. Compared to a joint model, SGM provides a 16.7×16.7$\times$16.7 × speedup in number of network updates and a 31.9×31.9$\times$31.9 × speedup in TFLOPs (see Fig. 5). Moreover, SGM requires 18×18$\times$18 × less training time than the joint model on the same hardware.
@@ 6.1 IID Continual Learning
Table 4: IID CL Results. Results after learning ImageNet-1K followed by Places365-LT over 5 rehearsal sessions. $\mu$ denotes average accuracy (%) over rehearsal sessions and $\alpha$ is final accuracy (%) on all 1365 classes. $\sigma$ stands for final accuracy (%) on ImageNet-1K only. #$\#P$ denotes trainable parameters in Millions. We report the average over $5$ data orderings with standard deviation ($\pm$).
@@ 6.2 Storage Constrained Offline Continual Learning
After being pre-trained on ImageNet-1K, a model sequentially learns 5 incremental batches of data (73 classes per batch) from Places365-Standard in offline CL with CIL ordering. Each rehearsal is compute-constrained where the model is updated over 1200 minibatches. Each minibatch consists of 256 samples where 50% are selected randomly from the current CL batch and 50% from data seen in earlier CL batches and ImageNet-1K. During learning a new batch, the model rehearses old data from storage which is bounded by a maximum number of samples e.g., 192K and 24K corresponding to $6.4\%$ and $0.8\%$ of the entire dataset (ImageNet and Places combined), respectively. As shown in Table 5, SGM improves each method’s performance in all criteria. When the storage is bounded by 24K samples, SGM improves final accuracy by absolute $11.24\%$ (DERpp) and $14.08\%$ (GDumb) and provides 2.4×2.4$\times$2.4 × and 3.1×3.1$\times$3.1 × more stability for DERpp and GDumb, respectively.
@@ 6.3 Storage Constrained Online Continual Learning
We also assess SGM’s efficacy in an online CL setting using a state-of-the-art online CL method, REMIND (Hayes et al., 2020). Using ConvNeXtV2-Femto, we conduct storage-constrained CL experiments with the CIL data ordering, where we combine SGM with REMIND. Since performing online CL experiments on large-scale datasets is computationally expensive, we choose a small dataset, CUB-200 as the 2nd dataset after ImageNet-1K. CUB-200 is more challenging than other small datasets (Lee et al., 2023). After being pre-trained on ImageNet-1K, a model learns CUB-200 in sample-by-sample manner, i.e., after receiving a new sample, it takes one SGD step using 51 samples (50 old samples and 1 new sample). Storage is constrained by a maximum number of samples, i.e., 80K and 20K, corresponding to $6.2\%$ and $1.5\%$ of the entire dataset (ImageNet and CUB combined), respectively. As shown in Table 6, SGM combined with REMIND outperforms standalone REMIND by large margins in all metrics. When storage is bounded by 20K samples, SGM improves REMIND’s final accuracy by absolute $8.64\%$ and reduces REMIND’s stability gap by a factor of 3.95×3.95$\times$3.95 ×.
@@ 6.4 Supporting Studies
Class-Balanced Rehearsal. Since our main experiments are based on conventional rehearsal without class balance, we also conduct experiments using class-balanced rehearsal. Compared to vanilla rehearsal, SGM reduces the stability gap by 7.3×7.3$\times$7.3 × when both use class-balanced rehearsal. SGM also achieves continual knowledge transfer (see Appendix D.2).
Supervised Pre-Training. In our main results, SGM reduces the stability gap using the self-supervised pre-trained ConvNeXtV2 model, and self-supervised models are known to perform better in CL (Gallardo et al., 2021). We also conduct experiments with the supervised pre-trained ConvNeXtV1 (Liu et al., 2022) to examine how SGM performs without a self-supervised backbone. We find that SGM is effective without the self-supervised backbone. Compared to vanilla rehearsal, SGM achieves 6×6$\times$6 × more stability, 3.9×3.9$\times$3.9 × more plasticity, and 35×35$\times$35 × more continual knowledge transfer when both use ConvNeXtV1 (see Appendix D.3).
Vision Transformers. We also study the behavior of SGM with ViTs, which now rival CNNs. Compared to vanilla rehearsal, SGM reduces the stability gap by 2.4×2.4$\times$2.4 × when both use a pre-trained ViT (see Appendix D.4).
@@ 7 Limitations & Future Work
Our work is built upon the assumption that a well-trained DNN is provided and performance on the original tasks must be maintained or improved. If this assumption is violated, as seen in some CL papers that do not start with pre-trained models, SGM’s dependence on LoRA would impair its utility. Although most of the components of SGM, such as weight initialization, dynamic soft targets, and OOCF, can be readily applied to DNNs trained from scratch, LoRA cannot be used because DNN layers without LoRA would remain frozen and unlearned. Applying SGM in this setting would require a variant of LoRA where LoRA could be used for all layers from the start of training such that the DNN rivaled a jointly trained model without LoRA. We used pre-trained ConvNeXtV1, ConvNeXtV2, and MobileViT DNNs, which are amenable to LoRA. Future work could investigate different LoRA variants and network adapters (Han et al., 2024) with SGM. It would also be interesting to study the efficacy of SGM for meeting the needs of CL for embedded devices Hayes & Kanan (2022).
Our study focused on image classification. Future work could explore SGM’s efficacy for regression, object detection (Acharya et al., 2020; 2019), reasoning Hayes & Kanan (2021), and semantic segmentation (Zhang et al., 2022). CL methods developed only on small datasets often do not scale to larger datasets such as ImageNet-1K (Zhou et al., 2023), so we studied a combination of ImageNet-1K and Places365. Due to computational limitations and a lack of readily available large-scale image classification datasets, we could not study CL for more than 1365 classes. In future work, assessing how well SGM scales with increasing dataset sizes and class counts would be valuable. We hope our work will inspire future research to find better approaches to prevent model decay due to concept drift.
@@ 8 Conclusion
Although there has been remarkable progress in mitigating catastrophic forgetting, almost all CL methods remain unsuited for practical applications such as addressing model decay (Verwimp et al., 2024; Harun et al., 2023a). To eliminate the need for frequent retraining of deployed DNNs, CL has to combat model decay, mitigate the stability gap, and be more computationally efficient than retraining from scratch in large-scale CL settings. Focusing on this scenario, we showed that SGM meets these criteria. With the growing energy usage of deep learning models (Luccioni et al., 2022; Patterson et al., 2021; Wu et al., 2022), reducing the need for retraining from scratch could significantly contribute to reducing carbon emissions associated with training DNNs. Our work aligns CL with these real-world challenges, and we hope it encourages the CL community to focus on them.
@@ Acknowledgments
This work was supported in part by NSF awards #1909696, #2326491, #2125362, and #2317706. The views and conclusions contained herein are those of the authors and should not be interpreted as representing any sponsor’s official policies or endorsements. We thank Jhair Gallardo, Junyu Chen, Shikhar Srivastava, and Robik Shrestha for their feedback and comments on the manuscript.
@@ Appendix A Implementation Details and Additional Results
  • Appendix G illustrates the process of updating soft targets.
  • Appendix H includes additional discussion.
@@ Appendix C Additional Implementation Details
Storage Constrained Offline CL with DERpp and GDumb. We describe settings used in Sec. 6.2 where we combine SGM with DERpp and GDumb while using identical settings e.g., the same ImageNet-1K pre-trained ConvNeXt V2 Femto network and the same optimizer settings. Each model pre-trained on ImageNet-1K learns Places365-Standard in 5 batches subsequently (73 categories per batch). Each rehearsal session performs a total of $1200$ iterations with $256$ samples per iteration. DERpp employs distillation and regularization along with rehearsal to prevent catastrophic forgetting. It regularizes loss on old samples and uses an additional distillation loss on logits of old samples for promoting consistency. We set coefficients $\alpha=0.1$ and $\beta=0.9$ for distillation and regularization respectively. GDumb randomly removes a sample from the largest class when the buffer reaches its maximum capacity and maintains a class-balanced buffer. For all methods, the buffer is bounded by a maximum number of samples ($80\%$ ImageNet-1K + $20\%$ Places365-Standard). DERpp, GDumb, and SGM use an initial learning rate of $1\times 10^{-3}$, $1\times 10^{-3}$, and $1.5\times 10^{-3}$, respectively, for batch size $256$. The joint model (upper bound) uses an initial learning rate of $10^{-2}$ and $12$K iterations with $256$ samples per iteration. We assess performance during rehearsal every 100 mini-batches to compute the metrics.